OpenStudy (anonymous):

(3x2 + 7x - 18) (x - 3)

OpenStudy (abb0t):

distribute: \[(x \times 3x^2)+(x \times 7x)-(x \times 18) - [(3 \times 3x^2) - (3 \times 7x) - (3 \times -18)\]

OpenStudy (anonymous):

so now i just multiply all them